# Intelligent Agricultural Greenhouse Control System Based on Internet of Things and Machine Learning

### v1

@@ Abstract
This study endeavors to conceptualize and execute a sophisticated agricultural greenhouse control system grounded in the amalgamation of the Internet of Things (IoT) and machine learning. Through meticulous monitoring of intrinsic environmental parameters within the greenhouse and the integration of machine learning algorithms, the conditions within the greenhouse are aptly modulated. The envisaged outcome is an enhancement in crop growth efficiency and yield, accompanied by a reduction in resource wastage. In the backdrop of escalating global population figures and the escalating exigencies of climate change, agriculture confronts unprecedented challenges. Conventional agricultural paradigms have proven inadequate in addressing the imperatives of food safety and production efficiency. Against this backdrop, greenhouse agriculture emerges as a viable solution, proffering a controlled milieu for crop cultivation to augment yields, refine quality, and diminish reliance on natural resources [b1]. Nevertheless, greenhouse agriculture contends with a gamut of challenges. Traditional greenhouse management strategies, often grounded in experiential knowledge and predefined rules, lack targeted personalized regulation, thereby resulting in resource inefficiencies. The exigencies of real-time monitoring and precise control of the greenhouse’s internal environment gain paramount importance with the burgeoning scale of agriculture. To redress this challenge, the study introduces IoT technology and machine learning algorithms into greenhouse agriculture, aspiring to institute an intelligent agricultural greenhouse control system conducive to augmenting the efficiency and sustainability of agricultural production.
@@
1 Introduction
In the formulation of the intelligent agricultural greenhouse control system, ethical considerations and system comprehensiveness demand meticulous attention. The real-time monitoring of greenhouse environmental parameters necessitates the collection and processing of copious sensitive data, thereby engendering privacy concerns for agricultural stakeholders, greenhouse equipment manufacturers, and other involved parties. In response, the study employs secure data transmission and storage mechanisms that encrypt sensitive information, ensuring exclusive access to critical data by authorized personnel. Additionally, a well-defined data use policy delineates ownership and usage rights, complemented by data desensitization measures to mitigate potential privacy risks.
@@ 1 Introduction
The study’s objective revolves around the design and implementation of an intelligent agricultural greenhouse control system predicated on the IoT and machine learning, with the overarching aim of enhancing crop growth efficiency and yield. By scrutinizing internal environmental parameters such as temperature, humidity, and light, coupled with the judicious application of machine learning algorithms, real-time and precise control of these parameters is sought. The anticipated outcome is an adaptive system enabling farmers to cater to diverse crop needs, mitigate resource wastage, and elevate production efficiency.
@@
Within agriculture, machine learning algorithms find widespread utility in forecasting, optimization, and decision support. Historical data analysis empowers the construction of models predicting crop growth under diverse meteorological conditions, thereby enabling precise regulation of growth environments. Furthermore, machine learning can dynamically adjust greenhouse conditions based on real-time data, aligning with distinct stages of crop growth to amplify production efficiency.
@@
A discernible focus of the study is the resolution of the comprehensive optimization challenge in greenhouse environments. The recognition of a research gap in the extant literature pertaining to the holistic optimization of greenhouse environmental parameters propels the study to center on comprehensive greenhouse environment optimization, including temperature, humidity, light, and other facets. The overarching objective is to fashion a system that comprehensively meets the diverse needs of crop growth across distinct stages.
@@
2 Greenhouse internal environmental monitoring system
@@
2.1 Sensor selection and layout: temperature, humidity, light and other parameters
@@
Sensor Layout
The efficacy of the monitoring system hinges on the meticulous arrangement of sensors. We evenly distribute temperature and humidity sensors throughout diverse areas of the greenhouse to obtain comprehensive environmental data [2-3]. The light sensor is strategically positioned to discern changes in light, enabling real-time monitoring of light conditions.
2.2 Data acquisition and transmission: real-time monitoring of the greenhouse internal environment
@@
Table 1: Instructions and Their Functions
@@
AT+CWMODE=1
Set the WiFi module to STA mode
AT+RST
Restart the WiFi module to take effect
AT+CWJAP=”xx”,”xxxxxxxx”
Connect to a WiFi hotspot with SSID ”xx” and password ”xxxxxxxx”
AT+CIPMUX=0
Enable single connection mode
AT+CIPSTART=”TCP”, ”192.168.1.XXX”, 8000
Establish a TCP connection to IP 192.168.1.XXX on port 8000
AT+CIPMODE=1
Enable transparent transmission mode
AT+CIPSEND
Begin data transmission
2.3 Data processing and storage: Cloud computing applications
@@
2.4 Data processing and storage: Cloud computing applications
Data processing
@@ 2.4 Data processing and storage: Cloud computing applications
Anomaly Detection and Repair During Data Processing:
@@ 1.1 Standard deviation method:
Application of machine learning algorithm
@@
3 Application of machine learning algorithms
3.1 Data analysis and model training: The use of historical data
@@
Feature engineering
Preceding model training, a crucial step involves feature engineering and the judicious selection of pertinent feature variables. Notably, features such as temperature, humidity, and illuminance over a temporal span serve as input variables, while future temporal points’ environmental parameters serve as output labels.
Model selection
Given the sequential and dynamic nature of greenhouse environmental changes, the adoption of a Recurrent Neural Network (RNN) for model training is deemed appropriate. RNN, a deep learning model tailored for sequential data, exhibits a fundamental structure as expressed below:
@@
The pivotal roles of data preprocessing and feature engineering in model performance underscore the importance of meticulous feature selection related to dynamic changes in the greenhouse environment, thereby augmenting the model’s generalization capabilities. Normalization and standardization of data ensure uniform feature value ranges, thereby facilitating the stable training of the model. These adaptive methodologies collectively aim to effectively surmount diverse challenges encountered in model training, thereby enhancing the performance and reliability of the intelligent agricultural greenhouse control system.
3.2 Prediction of greenhouse conditions: a model based on machine learning
@@
3.3 Overview of control strategies and algorithms of real-time regulation system:
@@
Figure 3: Measurement results using the smart agricultural system (Air Humidity, Soil Moisture (left)and Soil Temperatire, Air Temperature (right)
@@
Model Predictive Control (MPC)
@@
Fuzzy control system
@@
Integrated Control Policy
@@
Impact of Real-Time Regulation
@@
Table 2: Units for Magnetic Properties
Number
Air Temperature
Air Humidity %
Soil Temperature
Soil Moisture %
CO2 Concentration (PPM)
1
21.3 C
29.1
25.5
3000.1
441
2
21.6 C
29.1
25.6
3000.2
444
3
21.6 C
29.3
25.7
3000.2
449
4
21.7 C
30.0
25.8
2999.2
455
5
21.9 C
30.1
26.9
2999.1
459
6
21.9 C
30.1
27.1
3000.2
467
7
22.1 C
29.1
29.5
3001.2
471
8
22.2 C
29.2
27.8
3005.1
473
9
23.2 C
30.1
28.2
3010.1
482
10
23.5 C
30.0
28.9
3011.2
491
Table 3: The testing results of the Dedicated testing tools
Number
Air Temperature
Air Humidity %
Soil Temperature
Soil Moisture %
CO2 Concentration (PPM)
1
20.3 C
28.1
25.2
2999.1
439
2
20.5 C
28.1
25.2
2995.2
442
3
20.6 C
28.2
25.3
2991.2
448
4
20.5 C
29.0
25.5
2988.2
452
5
20.6 C
29.0
26.1
2995.1
458
6
20.7 C
29.0
26.5
2999.2
465
7
21.3 C
28.0
26.9
3000.2
469
8
21.8 C
29.0
27.1
3004.1
472
9
22.1 C
29.2
27.9
3009.1
480
10
22.5 C
29.3
28.2
3010.2
489
@@
4 Improvement of crop growth efficiency
4.1 Optimization of the growing environment: according to the needs of different crops
@@
Adjustment of growth environment parameters
@@
Individual regulation of different crops
@@
4.2 Improvement of resource utilization efficiency
@@
Energy consumption optimization
@@
Reduce fertilizer and water use
@@
5 Energy saving and environmental protection considerations
5.1 Optimization of energy consumption
@@
Energy consumption model
@@
5.2 Reduce fertilizer and water use
@@
Water resource utilization model
@@
Fertilization optimization algorithm
@@
6 Ensuring System Stability and Reliability
6.1 System Stability and Reliability
@@
System Status Monitoring
@@
7 Conclusion
7.1 Summarize the main findings of the study
@@
Effectiveness of intelligent greenhouse control system
@@
The key role of sensor technology and machine learning algorithms
@@
7.2 Outlook on the future development of intelligent agricultural greenhouse control system
Introducing a new generation of sensor technology
@@ 7.2 Outlook on the future development of intelligent agricultural greenhouse control system
Explore advanced deep learning algorithms
With the improvement of computing power, the application of advanced deep learning algorithms in greenhouse control systems will continue to be explored in the future. With more complex pattern recognition and learning capabilities, the system will better adapt to the growth characteristics of different crops and cope with complex and changing environmental conditions.

### v2

@@ Abstract
This study endeavors to conceptualize and execute a sophisticated agricultural greenhouse control system grounded in the amalgamation of the Internet of Things (IoT) and machine learning. Through meticulous monitoring of intrinsic environmental parameters within the greenhouse and the integration of machine learning algorithms, the conditions within the greenhouse are aptly modulated. The envisaged outcome is an enhancement in crop growth efficiency and yield, accompanied by a reduction in resource wastage. In the backdrop of escalating global population figures and the escalating exigencies of climate change, agriculture confronts unprecedented challenges. Conventional agricultural paradigms have proven inadequate in addressing the imperatives of food safety and production efficiency. Against this backdrop, greenhouse agriculture emerges as a viable solution, proffering a controlled milieu for crop cultivation to augment yields, refine quality, and diminish reliance on natural resources [1]. Nevertheless, greenhouse agriculture contends with a gamut of challenges. Traditional greenhouse management strategies, often grounded in experiential knowledge and predefined rules, lack targeted personalized regulation, thereby resulting in resource inefficiencies. The exigencies of real-time monitoring and precise control of the greenhouse’s internal environment gain paramount importance with the burgeoning scale of agriculture. To redress this challenge, the study introduces IoT technology and machine learning algorithms into greenhouse agriculture, aspiring to institute an intelligent agricultural greenhouse control system conducive to augmenting the efficiency and sustainability of agricultural production.
Index Terms:
@@
I Introduction
In the formulation of the intelligent agricultural greenhouse control system, ethical considerations and system comprehensiveness demand meticulous attention. The real-time monitoring of greenhouse environmental parameters necessitates the collection and processing of copious sensitive data, thereby engendering privacy concerns for agricultural stakeholders, greenhouse equipment manufacturers, and other involved parties[2]. In response, the study employs secure data transmission and storage mechanisms that encrypt sensitive information, ensuring exclusive access to critical data by authorized personnel. Additionally, a well-defined data use policy delineates ownership and usage rights, complemented by data desensitization measures to mitigate potential privacy risks.
@@ I Introduction
The study’s objective revolves around the design and implementation of an intelligent agricultural greenhouse control system predicated on the IoT and machine learning, with the overarching aim of enhancing crop growth efficiency and yield[3]. By scrutinizing internal environmental parameters such as temperature, humidity, and light, coupled with the judicious application of machine learning algorithms, real-time and precise control of these parameters is sought. The anticipated outcome is an adaptive system enabling farmers to cater to diverse crop needs, mitigate resource wastage, and elevate production efficiency.
@@
Within agriculture, machine learning algorithms find widespread utility in forecasting, optimization, and decision support. Historical data analysis empowers the construction of models predicting crop growth under diverse meteorological conditions, thereby enabling precise regulation of growth environments. Furthermore, machine learning can dynamically adjust greenhouse conditions based on real-time data, aligning with distinct stages of crop growth to amplify production efficiency[4].
@@
A discernible focus of the study is the resolution of the comprehensive optimization challenge in greenhouse environments. The recognition of a research gap in the extant literature pertaining to the holistic optimization of greenhouse environmental parameters propels the study to center on comprehensive greenhouse environment optimization, including temperature, humidity, light, and other facets[5]. The overarching objective is to fashion a system that comprehensively meets the diverse needs of crop growth across distinct stages.
@@
II Greenhouse internal environmental monitoring system
@@
II-A Sensor selection and layout: temperature, humidity, light and other parameters
@@ II-A1 Sensor Selection
II-A2 Sensor Layout
The efficacy of the monitoring system hinges on the meticulous arrangement of sensors. We evenly distribute temperature and humidity sensors throughout diverse areas of the greenhouse to obtain comprehensive environmental data [6]. The light sensor is strategically positioned to discern changes in light, enabling real-time monitoring of light conditions.
II-B Data acquisition and transmission: real-time monitoring of the greenhouse internal environment
@@ II-B2 Data Transfer
TABLE I: Instructions and Their Functions
@@
AT+CWMODE=1 Set the WiFi module to STA mode
AT+RST Restart the WiFi module to take effect
AT+CWJAP=”xx”,”xxxxxxxx” Connect to a WiFi hotspot with SSID ”xx” and password ”xxxxxxxx”
AT+CIPMUX=0 Enable single connection mode
AT+CIPSTART=”TCP”, ”192.168.1.XXX”, 8000 Establish a TCP connection to IP 192.168.1.XXX on port 8000
AT+CIPMODE=1 Enable transparent transmission mode
AT+CIPSEND Begin data transmission
II-C Data processing and storage: Cloud computing applications
@@
II-C1 Data processing
@@
Anomaly Detection and Repair During Data Processing
@@ 1.1 Standard deviation method:
II-C2 Application of machine learning algorithm
@@
III Application of machine learning algorithms
III-A Data analysis and model training: The use of historical data
@@ III-A1 Data Preparation
III-A2 Feature engineering
Preceding model training, a crucial step involves feature engineering and the judicious selection of pertinent feature variables. Notably, features such as temperature, humidity, and illuminance over a temporal span serve as input variables, while future temporal points’ environmental parameters serve as output labels[8].
III-A3 Model selection
Given the sequential and dynamic nature of greenhouse environmental changes, the adoption of a Recurrent Neural Network (RNN) for model training is deemed appropriate[9]. RNN, a deep learning model tailored for sequential data, exhibits a fundamental structure as expressed below:
@@ III-A3 Model selection
The pivotal roles of data preprocessing and feature engineering in model performance underscore the importance of meticulous feature selection related to dynamic changes in the greenhouse environment, thereby augmenting the model’s generalization capabilities[10]. Normalization and standardization of data ensure uniform feature value ranges, thereby facilitating the stable training of the model. These adaptive methodologies collectively aim to effectively surmount diverse challenges encountered in model training, thereby enhancing the performance and reliability of the intelligent agricultural greenhouse control system.
III-B Prediction of greenhouse conditions: a model based on machine learning
@@
III-C Overview of control strategies and algorithms of real-time regulation system:
@@
Figure 3: Measurement results using the smart agricultural system (Air Humidity, Soil Moisture (left)and Soil Temperatire, Air Temperature (right))
@@ III-C1 PID Controller
III-C2 Model Predictive Control (MPC)
@@
III-C3 Fuzzy control system
@@
III-C4 Integrated Control Policy
@@
III-C5 Impact of Real-Time Regulation
@@
TABLE II: Units for Magnetic Properties
Number Air Temperature Air Humidity % Soil Temperature Soil Moisture % CO2 Concentration (PPM)
1 21.3 C 29.1 25.5 3000.1 441
2 21.6 C 29.1 25.6 3000.2 444
3 21.6 C 29.3 25.7 3000.2 449
4 21.7 C 30.0 25.8 2999.2 455
5 21.9 C 30.1 26.9 2999.1 459
6 21.9 C 30.1 27.1 3000.2 467
7 22.1 C 29.1 29.5 3001.2 471
8 22.2 C 29.2 27.8 3005.1 473
9 23.2 C 30.1 28.2 3010.1 482
10 23.5 C 30.0 28.9 3011.2 491
TABLE III: The testing results of the Dedicated testing tools
Number Air Temperature Air Humidity % Soil Temperature Soil Moisture % CO2 Concentration (PPM)
1 20.3 C 28.1 25.2 2999.1 439
2 20.5 C 28.1 25.2 2995.2 442
3 20.6 C 28.2 25.3 2991.2 448
4 20.5 C 29.0 25.5 2988.2 452
5 20.6 C 29.0 26.1 2995.1 458
6 20.7 C 29.0 26.5 2999.2 465
7 21.3 C 28.0 26.9 3000.2 469
8 21.8 C 29.0 27.1 3004.1 472
9 22.1 C 29.2 27.9 3009.1 480
10 22.5 C 29.3 28.2 3010.2 489
@@
IV Improvement of crop growth efficiency
IV-A Optimization of the growing environment: according to the needs of different crops
@@
IV-A1 Adjustment of growth environment parameters
@@
IV-A2 Individual regulation of different crops
@@
IV-B Improvement of resource utilization efficiency
@@
IV-B1 Energy consumption optimization
@@
IV-B2 Reduce fertilizer and water use
@@
V Energy saving and environmental protection considerations
V-A Optimization of energy consumption
@@
V-A1 Energy consumption model
@@ V-A2 Control Policy
V-B Reduce fertilizer and water use
@@
V-B1 Water resource utilization model
@@
V-B2 Fertilization optimization algorithm
@@
VI Ensuring System Stability and Reliability
VI-A System Stability and Reliability
@@
VI-A1 System Status Monitoring
@@ VI-A2 Redundancy Design
VII Conclusion
VII-A Summarize the main findings of the study
@@
VII-A1 Effectiveness of intelligent greenhouse control system
@@
VII-A2 The key role of sensor technology and machine learning algorithms
@@
VII-B Outlook on the future development of intelligent agricultural greenhouse control system
VII-B1 Introducing a new generation of sensor technology
@@
VII-B2 Explore advanced deep learning algorithms
With the improvement of computing power, the application of advanced deep learning algorithms in greenhouse control systems will continue to be explored in the future. With more complex pattern recognition and learning capabilities, the system will better adapt to the growth characteristics of different crops and cope with complex and changing environmental conditions[12].